OpenStudy (anonymous):

I WILL FAN AND MEDAL Use the formulas to generate a Pythagorean Triple with x = 5 and y = 2. The three side lengths, from smallest to largest are: ___,___ and ___

OpenStudy (swaqqout_kid1):

is there a graph

OpenStudy (anonymous):

nope @SwaqqOut_KID1

OpenStudy (swaqqout_kid1):

o k

OpenStudy (swaqqout_kid1):

i am goin 2 get som help ight

OpenStudy (swaqqout_kid1):

@sangya21 @sammixboo

OpenStudy (swaqqout_kid1):

confused :(

OpenStudy (anonymous):

pythagoras theorem is x^2 +y^2 = z^2 (5)^2 +(2^2) = z^2 25+4 = z^2 29 = z^2 sqrt(29) = z arrange it in smallest to largest

OpenStudy (swaqqout_kid1):

y thank you @sangya21

OpenStudy (anonymous):

I hope it helps. so anytime :)

OpenStudy (anonymous):

@sangya21 Can you help. I still dont understand

OpenStudy (anonymous):

A Pythagorean triple is a set of three integers a, b and c that specify the lengths of a right triangle - that is c^2 = a^2 + b^2. a = x, b = y c^2 = x^2 + y^2